The maximum distance between the Earth and Mars is 378,000,000 km and the minimum distance is 168,000,000 km. The maximum amount of time for a radio signal to travel from Mars to Earth is ____ seconds.
Answer #1

For the maximum time, the distnace between mars and earth should be maximum

Maximum distnace=378000000 km=3.78*10^11 m

Speed of radio signal=3*10^8 m/s

Time taken by radio signal=distance/speed=3.78*10^11/(3*10^8)=1.26*10³ seconds=1260 seconds